OpenStudy (anonymous):

helppppp what is the probability of getting heads when flipping a coin and getting a number greater than or equal to 3 when rolling a single die

OpenStudy (xapproachesinfinity):

So P(heads and #>=3) this are two independent events first let's see what is the P(of head)=1/2 1 in two chances right?

OpenStudy (xapproachesinfinity):

we have 6 side in the die so total chances numbers that are greater or equal to 3 are 3,4,5,6 so 4/6 4 in 6 chances. are you following?

OpenStudy (anonymous):

yeah and rolligna die is 1/6 ?

OpenStudy (anonymous):

oh

OpenStudy (xapproachesinfinity):

1/6 chance for each side that is

OpenStudy (xapproachesinfinity):

they are asking event that a number greater than or equal to 3 to occurs in this experiment do you get it?

OpenStudy (xapproachesinfinity):

Now you can use this P(A and B)=P(A).P(B) okay?

OpenStudy (anonymous):

yes i think

OpenStudy (xapproachesinfinity):

A event that we get head B event that we get a number greater than or equal to 3

OpenStudy (xapproachesinfinity):

we have already got P(A) and P(B) just multiply

OpenStudy (xapproachesinfinity):

can you do it?

OpenStudy (anonymous):

yeah but i dont see where it says what p(a) and p(b) are is it the fraction?

OpenStudy (xapproachesinfinity):

P(head)=1/2 i did it above P(#>=3)=4/6=2/3 also i did it above P(head and #>=3)=P(Head)P(#>=3)

OpenStudy (xapproachesinfinity):

i was just linking the P(A and B)=P(A).P(B) that you studied with this

OpenStudy (anonymous):

is the answer 1/3

OpenStudy (xapproachesinfinity):

Correct!

OpenStudy (anonymous):

ok thanks!!!!

OpenStudy (xapproachesinfinity):

You're welcome! you are good to go with this problems haha

OpenStudy (xapproachesinfinity):

just apply the same thing in case of independent events